Help!
American Products plans to lay off 10% of it's employees in its aerospace division, and 15% of its employees in its agriculture division. If altogether 12% of the 3000 employees in these two divisions will be laid off, then how many employees are in each division?

Answer by stanbon(48558)

American Products plans to lay off 10% of it's employees in its aerospace division, and 15% of its employees in its agriculture division. If altogether 12% of the 3000 employees in these two divisions will be laid off, then how many employees are in each division?
There are 3000 employees in the two divisions.
Let number in aerospace be "x"
Then number of agriculture is "3000-x"
Number laid off in aerospace is 10%(x) = 0.10x
Number laid off in agriculture is 15%(3000-x)=450-0.15x
Number laidoff overall is 12%(3000)=360
EQUATION:
laid off + laid off= 360
0.10x + 450-0.15x=360
-0.05x=-90
x=1800 (number in aerospace)
3000-x=1200 (number in agriculture)
Cheers,
Stan H.